What is |6+3a|= 15 i need a solved
Answer by stanbon(75887)

What is |6+3a|= 15
-----
That "thing" inside the absolute value symbols
must be either +15 or -15.
--------------------------
Solve:
6+3a = 15 or 6+3a = -15
---
3a = 9 or 3a = -21
---
a = 3 or a = -7
===========================
Cheers,
Stan H.
===========================